Grade 12Mechanics

A TRAIN OF MASS “M” IS MOVING WITH UNIFORM VELOCITY.SUDDENLY THE LAST BOGGY OF THE TRAIN OF MASS “m” DETTACHES FROM IT.THE TRAIN STOPS GOING A DISTANCE “L”.FIND THE DISTANCE BETWEEN THE TRAIN AND THE BOGGY WHEN BOTH OF THEM COME TO REST.(the force opposing the motion of the train is proportional to its mass and the tension of the engine is constant.)

sayandeepa raha
11 Years agoGrade 12

1 Answer

kurapati bhanuraja
11 Years ago
F is directly proportional to mass.i.e, intially F=k(M+m). when the boggy is detached individual forces acting on them are.F1=kM,F2=km.let the velocity during their seperation be ‘v’.this is the intial velocity during their datachment and after detachment they retard with accelerations a1,a2.given the distance travelled by M is L 02-v2=-2a1L.and 02-v2=-2a2s.hence from the relation of a1,a2 u can bring the relation among ‘s’ and L. and distance between them is ‘L-s’.i think the problem can also be solved by centre of mass concept